OpenStudy (anonymous):

sin^2+cos^2+Tan^2=sec^2

OpenStudy (solomonzelman):

\(\LARGE \color{blue}{ \bf sin^2x+cos^2x+tan^2x=sec^2x }\) \(\LARGE \color{blue}{ \bf (1-cos^2x)+cos^2x+tan^2x=sec^2x }\) \(\LARGE \color{blue}{ \bf 1+tan^2x=sec^2x }\) identity verified.

OpenStudy (solomonzelman):

I substituted 1-cos²x for sin²x and then simplified it. So I am now facing one of the Pythagorean identities.

OpenStudy (solomonzelman):

More help needed ?